# PaperRegister: Boosting Flexible-grained Paper Search via Hierarchical Register Indexing

### v1

@@ Conclusion
Appendix B
This part is about detailed process and instructions for the fine-grained content extracting.
Learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction (Edge et al. 2024; Li et al. 2025c, 2024b, 2024c), PaperRegister uses a large language model as the extracting module. PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. In addition, PaperRegister also retrieve relevant paper text after the extracting as supplement to improve the register content. The instruction used for LLM is ==Instruction== You are an content extraction expert, particularly skilled at extracting structured records from academic papers. Below, I need you to perform extraction based on a given schema. Please note the following: 1. Your extraction does not need to preserve the original text verbatim. You may paraphrase or summarize the content to make the extracted content more comprehensive and fluent. 2. Not all field names in the schema will have corresponding content in the paper. If you cannot find a precise match in the paper, leave that field empty. 3. Your final output must strictly adhere to the original schema in JSON format, starting with ’“‘json’ and ending with ’“‘’. ==Schema== schema ==Paper== paper
Appendix C
This part is about detailed process and instructions for the bottom-up content extracting.
@@
Appendix D
This part is about detailed statistical information of dataset for the experiments.
@@
We first use LitSearch, mainly containing coarse-grained queries. And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training. Due to limited computational resources, we are unable to conduct experiments based on an ultra-large-scale paper corpus. For LitSearch, we extract 3,400 papers as the corpus, using the original 597 paper search queries. For Flexible-grained Search, we collect 4,200 papers as the corpus. The numbers of queries included in F.g.Search-1, F.g.Search-2, and F.g.Search-3 are 1,922, 1,922, and 1,800, respectively.

### v2

@@ Appendix A Detailed illustration for hierarchical register schema
Appendix B Detailed process and instructions for fine-grained content extracting
Learning from several widely recognized works, in which large language models are proven to possess reliable capabilities for content extraction edge2024local; li2025deepsolution; li2024structrag; li2024meta, PaperRegister uses a large language model as the extracting module. PaperRegister takes the node name and the text-formatted paper as input, uses instructions to guide in outputting corresponding content for the information node, and leaves it blank if the paper does not include corresponding content. In addition, PaperRegister also retrieve relevant paper text after the extracting as supplement to improve the register content. The instruction used for LLM is ==Instruction== You are an content extraction expert, particularly skilled at extracting structured records from academic papers. Below, I need you to perform extraction based on a given schema. Please note the following: 1. Your extraction does not need to preserve the original text verbatim. You may paraphrase or summarize the content to make the extracted content more comprehensive and fluent. 2. Not all field names in the schema will have corresponding content in the paper. If you cannot find a precise match in the paper, leave that field empty. 3. Your final output must strictly adhere to the original schema in JSON format, starting with ’“‘json’ and ending with ’“‘’. ==Schema== schema ==Paper== paper
Dataset Name # Corpus # Query
LitSearch 3400 597
F.g.Search-1 4200 1,922
F.g.Search-2 4200 1,922
F.g.Search-3 4200 1,800
Table 6: Statistics of test data in experiments.
Appendix C Detailed process and instructions for bottom-up content extracting
@@
Appendix D Detailed statistical information of dataset for experiments
@@
We first use LitSearch, mainly containing coarse-grained queries. And then we build a new dataset, Flexible-grained Search, covering queries across various granularity and including data for test, development, and training. Due to limited computational resources, we are unable to conduct experiments based on an ultra-large-scale paper corpus. For LitSearch, we extract 3,400 papers as the corpus, using the original 597 paper search queries. For Flexible-grained Search, we collect 4,200 papers as the corpus. The numbers of queries included in F.g.Search-1, F.g.Search-2, and F.g.Search-3 are 1,922, 1,922, and 1,800, respectively. The statics is in Table 6.
Appendix E Experimental results by more metrics
To provide a more comprehensive illustration of the advancements of PaperRegister over baseline methods, we present experimental results in the Table 789 across six metrics: Precision@5, Precision@10, MAP@5, MAP@10, NDCG@5, and NDCG@10. The data consistently support the conclusions drawn in our main experimental tables, further confirming that PaperRegister is an effective system for addressing flexible-grained paper search and the advantages of PaperRegister become more pronounced when handling finer-grained queries.
Figure 6: The first kind of hierarchical register schema (algorithm innovation).
Figure 7: The second kind of hierarchical register schema (benchmark construction).
Figure 8: The third kind of hierarchical register schema (mechanism exploration).
Figure 9: The fourth kind of hierarchical register schema (survey and review).
Figure 10: The fifth kind of hierarchical register schema (theory proof).
Method BM25-based Paper Search DPR-based Paper Search
P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10 P@5 P@10
Direct Matching
Title 11.4 6.3 8.4 4.7 7.2 4.2 6.1 3.5 14.0 7.8 10.4 5.9 9.2 5.2 8.1 4.7
Abstract 14.1 7.6 12.7 6.9 11.5 6.3 10.8 6.0 16.4 8.8 13.9 7.4 12.4 6.8 11.6 6.3
Total Paper 13.5 7.3 15.9 8.2 16.3 8.4 16.8 8.6 16.0 8.8 14.6 7.9 13.8 7.4 13.2 7.1
Query Enhancing
Rewriting ma2023query 13.0 7.4 11.7 6.4 10.8 6.0 10.2 5.6 16.0 8.8 13.4 7.3 12.1 6.5 11.0 6.1
HyDE gao2023precise 14.6 8.0 12.1 6.6 10.9 6.0 10.3 5.8 16.2 8.9 13.4 7.2 11.8 6.6 11.3 6.2
CSQE lei2024corpus 14.4 7.8 11.8 6.4 10.8 5.9 10.3 5.6 16.5 8.7 13.2 7.1 12.0 6.5 11.0 6.1
Multi-field Indexing
$\text{Chunk}_{avg}$ li2025multifield 12.2 7.2 13.6 7.4 13.3 7.2 13.2 7.2 10.5 6.2 10.2 5.8 9.3 5.3 9.3 5.2
$\text{Chunk}_{max}$ li2025multifield 14.3 7.9 16.0 8.3 16.3 8.5 17.1 8.8 15.2 8.5 16.0 8.3 15.3 8.1 15.8 8.2
$\text{Paragraph}_{avg}$ li2025multifield 6.2 4.1 11.8 6.7 12.0 6.7 11.6 6.6 4.3 2.4 4.7 3.1 4.6 3.0 4.5 2.9
$\text{Paragraph}_{max}$ li2025multifield 13.5 7.5 14.8 7.9 15.1 8.0 16.2 8.4 16.7 9.0 15.8 8.3 15.3 8.1 15.8 8.2
Hierarchical Register Indexing
PaperRegister (Ours) 14.6 8.0 17.9 9.1 17.6 8.9 17.5 8.9 17.1 9.3 16.8 8.7 16.0 8.3 16.2 8.3
Table 7: Precision@5 and Precision@10 results on paper search across various granularity.
Method BM25-based Paper Search DPR-based Paper Search
M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10 M@5 M@10
Direct Matching
Title 42.2 42.9 33.4 34.1 28.8 29.6 23.5 24.2 54.9 55.9 41.3 42.2 36.3 37.2 32.2 33.0
Abstract 52.4 53.0 54.9 55.6 49.4 50.1 45.8 46.5 65.1 65.9 59.8 60.5 51.9 52.7 47.2 47.8
Total Paper 51.7 52.4 74.3 74.6 76.1 76.4 79.1 79.3 60.5 61.6 63.2 64.0 59.2 59.9 55.6 56.3
Query Enhancing
Rewriting ma2023query 47.2 48.3 48.7 49.5 44.8 45.5 41.2 41.9 62.3 63.4 56.7 57.4 50.1 50.8 44.8 45.6
HyDE gao2023precise 54.1 55.1 50.1 50.9 44.6 45.3 41.7 42.6 64.3 65.3 56.4 57.1 48.7 49.6 45.1 45.9
CSQE lei2024corpus 54.1 54.8 41.2 42.0 37.8 38.4 36.7 37.4 65.2 65.8 54.1 54.9 48.4 49.1 44.4 45.1
Multi-field Indexing
$\text{Chunk}_{avg}$ li2025multifield 40.6 42.0 54.4 55.2 54.9 55.6 54.4 55.2 34.9 36.1 39.4 40.4 36.6 37.5 35.1 35.9
$\text{Chunk}_{max}$ li2025multifield 56.1 57.1 73.6 74.0 75.3 75.8 79.5 79.8 62.0 63.0 71.7 72.2 68.4 68.9 69.4 69.9
$\text{Paragraph}_{avg}$ li2025multifield 19.0 20.2 44.9 46.0 46.2 47.1 44.8 45.9 15.0 15.3 15.9 16.8 15.7 16.6 15.1 15.9
$\text{Paragraph}_{max}$ li2025multifield 50.7 51.7 66.4 67.0 67.2 67.8 73.2 73.6 66.1 67.0 70.8 71.3 68.5 69.1 68.7 69.1
Hierarchical Register Indexing
PaperRegister (Ours) 57.4 58.3 86.2 86.4 84.9 85.0 85.3 85.5 70.0 70.9 77.4 77.8 73.7 74.0 72.9 73.1
Table 8: MAP@5 and MAP@10 results on paper search across various granularity.
Method BM25-based Paper Search DPR-based Paper Search
N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10 N@5 N@10
Direct Matching
Title 45.4 47.1 35.6 37.3 30.6 32.5 25.2 26.8 58.2 60.3 44.0 46.2 38.7 40.7 34.3 36.4
Abstract 56.3 57.7 57.1 58.8 51.4 53.1 47.9 49.6 68.5 70.3 62.2 63.8 54.5 56.4 50.0 51.5
Total Paper 55.0 56.6 75.6 76.5 77.5 78.1 80.3 81.0 64.4 66.8 65.7 67.5 61.6 63.3 58.2 59.8
Query Enhancing
Rewriting ma2023query 51.2 53.7 51.1 53.0 47.1 48.9 43.6 45.4 66.0 68.4 59.3 61.1 52.7 54.3 47.4 49.3
HyDE gao2023precise 58.1 60.3 52.7 54.6 47.0 48.9 44.2 46.1 67.6 69.9 59.0 60.8 51.3 53.5 47.9 49.8
CSQE lei2024corpus 58.0 59.7 45.7 47.4 41.9 43.4 40.4 41.9 68.6 69.8 57.1 58.8 51.4 53.0 47.1 48.9
Multi-field Indexing
$\text{Chunk}_{avg}$ li2025multifield 45.2 48.4 57.8 59.8 57.8 59.5 57.3 59.1 38.8 41.8 42.3 44.7 39.1 41.3 37.9 39.8
$\text{Chunk}_{max}$ li2025multifield 59.2 61.7 75.2 76.2 76.8 77.9 81.0 81.7 64.9 67.3 73.8 74.8 70.5 71.7 71.8 72.8
$\text{Paragraph}_{avg}$ li2025multifield 21.8 24.7 48.4 50.9 49.6 51.7 48.2 50.8 16.4 17.2 17.7 20.0 17.5 19.7 16.9 19.0
$\text{Paragraph}_{max}$ li2025multifield 54.4 56.6 68.2 69.9 69.3 70.7 75.1 76.1 69.7 71.5 72.9 74.1 70.5 72.1 71.2 72.2
Hierarchical Register Indexing
PaperRegister (Ours) 60.7 62.9 87.1 87.5 85.7 86.0 85.9 86.3 73.0 75.0 79.1 80.1 75.2 76.1 74.9 75.5
Table 9: NDCG@5 and NDCG@10 results on paper search across various granularity.